# Computation method for perfectoid purity and perfectoid BCM-regularity

### v4

@@ Appendix A Construction of functorial test perfectoids
Next, let us show that $T(R)$ is a test perfectoid over $R$. Let $B$ be a perfectoid $R$-algebra, then by Andre’s flatness lemma [BS]*Theorem 7.14, we may assume that $B$ contains a compatible system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. Let $(\widetilde{B},(d^{\prime}))$ is a perfect prism such that $\widetilde{B}/(d^{\prime})\simeq B$. We consider the map $A_{\infty}=\mathbb{Z}_{(p)}[X^{1/p^{\infty}}]^{\wedge(p,d)}\to B$ by $X^{1/p^{e}}\mapsto p^{1/p^{e}}$, then it induces the homomorphism of perfectoids $A_{\infty}/(d)\to B$, and in particular, we obtain the homomorphism of perfect prisms $(A_{\infty},(d))\to(\widetilde{B},(d^{\prime}))$ by [BS]*Theorem 3.10. For each element $f\in R$, we take a lift $\widetilde{f}\in\widetilde{B}$ of the image of $f$ by $R\to B$, then we can define the homomorphism of perfect prisms $(A\{R\}_{\infty},(d))\to(\widetilde{B},(d^{\prime}))$ defined by $X_{f}\mapsto\widetilde{f}$, thus we obtain the homomorphism $Z\{R\}_{\infty}\to B$ and the commutative diagram
@@
Finally, let us show that $T(R)$ is $p$-torsion free if $R$ is $p$-torsion free. Since the endomorphism $\phi\colon A\{R\}\to A\{R\}$ is $(p,d)$-completely faithfully flat, so is $A\{R\}\to A\{R\}_{\infty}$. Since $A[R]\to A\{R\}$ is $(p,d)$-completely faithfully flat, so is $A[R]\to A\{R\}_{\infty}$. Therefore, we obtain that $\mathbb{Z}_{(p)}[R]\to Z\{R\}$ is $p$-completely faithfully flat, and in particular, so is $R\to Z\{R\}\otimes_{Z_{(}p)[R]}R$. Since $R$ is $p$-torsion free, so is $Z\{R\}\otimes_{Z_{(p)}[R]}R$. By [MSTWW]*Lemma A.2, we obtain that $T(R)$ is $p$-torsion free, as desired. ∎
@@ Appendix B Test ideal of graded ring
Since $S$ is quasi-Gorenstein, we obtain $\omega_{S}\simeq S(\nu)$ as graded $S$-modules, then $\nu<0$ by $a(S)<0$. By the isomorphism, the trace map $F_{*}\omega_{S}^{(p)}\to\omega_{S}$ is corresponding to $F_{*}(S(\nu))^{(p)}\to S(\nu)$, which is denoted by $u$. By the Matlis duality, it is enough to show that $\tau(S)=\bigcap_{e\in\mathbb{Z}_{\geq 0}}\mathrm{Im}(u^{e}\colon F^{e}_{*}S\to S)$, where $\tau(S)$ is the test ideal of $S$. First, let us show the following claim.
@@
Next, let us show the following claim.
@@
Next, let us show the following claim:
@@
Finally, let us show the assertion. We take rational numbers $01>t>\varepsilon>$ 0, then we have
@@
Appendix C Proof of Fedder type criterion
@@
We use Notation C.1. Let $n$ be a positive integer. Then $R$ is $n$-quasi-$F$-split if and only if there exists an $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that
@@
and in particular, $\psi$ satisfies the conditions (a) and (b). Next, we assume that there exists an $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that $\psi$ satisfies the conditions (a) and (b). Thus, the homomorphism $\psi$ induces the $\overline{R}$-module homomorphism $\overline{\psi}\colon Q_{R,n}\to\overline{R}$ such that $\overline{\psi}(1)$ is a unit. Replacing $\overline{\psi}$ by $\overline{\psi}(1)^{-1}\cdot\overline{\psi}$, we obtain a splitting of $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$, as desired. ∎
@@
We use Notation C.1. Let $n$ be an integer with $n\geq 2$. We take $\varphi\in\operatorname{Hom}_{\overline{A}}(Q_{A,n},\overline{A})$ and $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$. By the isomorphism $Q_{A,n}\simeq F_{*}Q_{A,n-1}\oplus F_{*}\overline{A}$ in the proof of Theorem 3.7, we obtain the homomorphism $\psi\in\operatorname{Hom}_{\overline{A}}(F_{*}Q_{A,n-1},\overline{A})$ and $h\in\overline{A}$ such that
@@
for every $\alpha=(a_{0},\ldots,a_{n-1})\in Q_{A,n}$. Furthermore, we assume $\varphi$ satisfies the condition (b) in Lemma C.2.
@@
Let us show (1) by induction on $n\geq 2$. We note that $h_{n}=g_{n}$ satisfies the desired condition for $s=n$. Since $\psi$ coincides with the decomposition
@@
and $\varphi$ satisfies the condition (b) in Lemma C.2, the homomorphism $\psi$ induces the $\overline{R}$-module homomorphism $\psi_{R}\colon F_{*}Q_{R,n-1}\to\overline{R}$. Since $\overline{R}$ is Gorenstein, the homomorphism $\psi_{R}$ factors through the homomorphism induced by $u(F_{*}(f^{p-1}\cdot-))\colon F_{*}\overline{A}\to\overline{A}$. Therefore, there exists $\varphi^{\prime}\in\operatorname{Hom}_{\overline{A}}(Q_{A,n-1},\overline{A})$ such that $\psi$ and $u(F_{*}(f^{p-1}\cdot-))\circ\varphi^{\prime}$ induce same homomorphism $Q_{R,n}\to\overline{R}$ and $\varphi^{\prime}$ satisfies the condition (b) in Lemma C.2. We take $g^{\prime}_{1},\ldots,g^{\prime}_{n-1}$ such that $\varphi^{\prime}=\varphi_{(g^{\prime}_{1},\ldots,g^{\prime}_{n-1})}$. By the isomorphism
@@
Next, let us show assertion (2). We take $r\geq j\geq 1$ and $a\in A$. Since $\varphi$ satisfies the condition (b), we obtain
@@
where we define $\Delta_{0}(af_{j})=af_{j}$. For $n-s\geq 3$ or $p\neq 2$, we have $g_{s}\Delta_{n-s}(af_{j})\in I^{[p^{n+1-s}]}$ by 3.6 and the assertion (1). Furthermore we have
@@
for every $n\geq 2$. Furthermore, $I^{\prime}_{n}\subseteq I^{\prime}_{n+1}$ for every $n\geq 1$ follows from the same argument. Let us show assertion (1). First, we consider the following claim.
@@
Let $g\in A$ and $n\geq 1$. Then $g\in I_{n}$ if and only if there exists $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2 and $g_{n}\equiv g\ \textrm{mod}\ I_{n-1}$, where if $n=1$, we regard $I_{0}$ as $(0)$.
@@
Let us show assertion by induction on $n$. For $n=1$, since $I_{1}=(I^{[p]}\colon I)$, $g\in I_{1}$ if and only if $u(F_{*}(g\cdot-))$ satisfies the condition (b) in Lemma C.2, as desired. Next, we assume $n\geq 2$. We take $g\in I_{n}$, then there exists $g^{\prime}\in I_{n-1}$ such that $g-u(F_{*}(g^{\prime}\Delta_{1}(f^{p-1})))\in I_{1}$. By the induction hypothesis, there exist $g^{\prime}_{1},\ldots,g^{\prime}_{n-1}$ such that $\varphi_{(g_{1}^{\prime},\ldots,g^{\prime}_{n-1})}$ satisfies the condition (b) in Lemma C.2, $g^{\prime}_{n-1}\equiv g^{\prime}\ \textrm{mod}\ I_{n-2}$, then by Lemma C.3, there exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g^{\prime}_{s}\equiv f^{p^{n-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n-s}]}$ for each $n-1\geq s\geq 1$. We define $\varphi^{\prime}$ by
@@
and $u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in I_{2}$, we obtain $g_{n}\equiv g\ \textrm{mod}\ I_{n}$. Thus, it is enough to show that $\varphi$ satisfies the condition (b) in Lemma C.2. We take $g_{1},\ldots,g_{n-1}$ such that $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$, then we have
@@
for $\alpha\in W_{n-1}((f_{1},\ldots,f_{r}))$, we obtain that $\varphi$ satisfies the condition (b) in Lemma C.2, as desired.
Next, let us show the converse direction, thus we take $g_{1},\ldots,g_{n}\in\overline{A}$ such that the corresponding homomorphism $\varphi:=\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2, then it is enough to show that $g_{n}\in I_{n}$. For $n=1$, we have $g_{1}\in(I^{[p]}\colon I))=I_{1}$, as desired, thus we assume $n\geq 2$. By Lemma C.3, there exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g_{s}\equiv f^{p^{n+1-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n+1-s}]}$, and in particular, $h_{n}\equiv g_{n}\ \textrm{mod}\ I^{[p]}$. By the proof of Lemma C.3 and the induction hypothesis, we obtain that $h_{n-1}\in I_{n-1}$, thus we have $u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))\in I_{n}$. Furthermore, by the construction of $I_{2}$, we have
@@
Finally, let us show the assertion. We have
@@
Next, let us show assertion (2). We define the sequence of ideals $\{I^{e}_{n}\}$ by $I^{e}_{1}:=f^{p-1}u^{e}(F^{e}_{*}f^{p^{e}-1}\overline{A})+I^{[p]}$ and
@@
Let $g\in A$ and $n,e\geq 1$. Then $g\in I^{e}_{n}$ if and only if there exists $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2, there exists $h\in f^{p^{e+n}-1}\overline{A}$ with $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, and $g_{n}\equiv g\ \textrm{mod}\ I^{e}_{n-1}$, where if $n=1$, we regard $I^{e}_{0}$ as $(0)$.
@@
For $n=1$, it follows from definition and the fact that the condition (b) in Lemma C.2 is equivalent to $g_{1}\in f^{p-1}\overline{A}+I^{[p]}$. We assume $n\geq 2$. We take $g\in I^{e}_{n}$, then there exists $g^{\prime}\in I^{e}_{n-1}$ such that $g-u(F_{*}(g^{\prime}\Delta_{1}(f^{p-1})))\in I_{1}$. By induction hypothesis, there exist $g_{1}^{\prime},\ldots,g^{\prime}_{n-1}$ such that $\varphi_{(g^{\prime}_{1},\ldots,g^{\prime}_{n-1})}$ satisfies the condition (b) in Lemma C.2, we have $u^{e}(F^{e}_{*}h^{\prime})=g^{\prime}_{1}$ for some $h^{\prime}\in f^{p^{e+n-1}-1}\overline{A}$, and $g^{\prime}_{n-1}\equiv g^{\prime}\ \textrm{mod}\ I^{e}_{n-2}$. By the proof of Claim C.4, if we put $g_{s}:=f^{p^{n-s}(p-1)}g^{\prime}_{s}$ for every $n-1\geq s\geq 1$ and
@@
then $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) Lemma C.2 and $g\equiv g_{n}\ \textrm{mod}\ I^{\prime}_{n}$. We note that $I_{2}\subseteq I^{e}_{n}$ if $n\geq 3$. Furthermore, we have
@@
Next, let us show the converse direction, thus we take $g_{1},\ldots,g_{n}\in\overline{A}$ such that the corresponding homomorphism $\varphi:=\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2 and there exists $h\in f^{p^{e+n}-1}\overline{A}$ such that $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, then it is enough to show that $g_{n}\in I^{e}_{n}$. For $n=1$, it follows from the definition of $I^{e}_{1}$. For $n\geq 2$, we take $h_{1},\ldots,h_{n}$ as in Lemma C.3, then the assertion follows from $h_{n-1}\in I^{e}_{n-1}$ by the induction hypothesis and
@@
for every $a\in\overline{A}$, and in particular, we have $u^{e}(F^{e}_{*}h)=g_{1}$. By Claim C.5, we have $g_{n}\in I^{e}_{n}=I^{\prime}_{n}$. Since $\varphi(1)=u(F_{*}g_{n})\notin\mathfrak{m}$, we have $I^{\prime}_{n}\nsubseteq\mathfrak{m}^{[p]}$, as desired. On the other hand, if $I^{e}_{n}=I^{\prime}_{n}\nsubseteq\mathfrak{m}^{[p]}$, then there exists $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$ such that $\varphi$ satisfies the condition (b) in Lemma C.2, there exists $h\in f^{p^{e+n}-1}\overline{A}$ such that $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, and $u(F_{*}g_{n})\notin\mathfrak{m}$ by Claim C.5. Then $(\varphi,u^{e+n}(F^{e+n}_{*}h\cdot-))\colon Q_{A,n}\oplus F^{e+n}_{*}\overline{A}\to\overline{A}$ induces the homomorphism $Q_{R,n,e}\to\overline{R}$ and $(\varphi,u(F^{e}_{*}h\cdot-))(1)$ is a unit, as desired. ∎

### v5

@@ Appendix A Construction of functorial test perfectoids
Next, we prove that $T(R)$ is a test perfectoid over $R$. Let $B$ be a perfectoid $R$-algebra, then by Andre’s flatness lemma [BS]*Theorem 7.14, we may assume that $B$ contains a compatible system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. Let $(\widetilde{B},(d^{\prime}))$ is a perfect prism such that $\widetilde{B}/(d^{\prime})\simeq B$. We consider the map $A_{\infty}=\mathbb{Z}_{(p)}[X^{1/p^{\infty}}]^{\wedge(p,d)}\to B$ by $X^{1/p^{e}}\mapsto p^{1/p^{e}}$, then it induces the homomorphism of perfectoids $A_{\infty}/(d)\to B$, and in particular, we obtain the homomorphism of perfect prisms $(A_{\infty},(d))\to(\widetilde{B},(d^{\prime}))$ by [BS]*Theorem 3.10. For each element $f\in R$, we take a lift $\widetilde{f}\in\widetilde{B}$ of the image of $f$ by $R\to B$, then we can define the homomorphism of perfect prisms $(A\{R\}_{\infty},(d))\to(\widetilde{B},(d^{\prime}))$ defined by $X_{f}\mapsto\widetilde{f}$, thus we obtain the homomorphism $Z\{R\}_{\infty}\to B$ and the commutative diagram
@@
Finally, we prove that $T(R)$ is $p$-torsion free if $R$ is $p$-torsion free. Since the endomorphism $\phi\colon A\{R\}\to A\{R\}$ is $(p,d)$-completely faithfully flat, so is $A\{R\}\to A\{R\}_{\infty}$. Since $A[R]\to A\{R\}$ is $(p,d)$-completely faithfully flat, so is $A[R]\to A\{R\}_{\infty}$. Therefore, we obtain that $\mathbb{Z}_{(p)}[R]\to Z\{R\}$ is $p$-completely faithfully flat, and in particular, so is $R\to Z\{R\}\otimes_{Z_{(}p)[R]}R$. Since $R$ is $p$-torsion free, so is $Z\{R\}\otimes_{Z_{(p)}[R]}R$. By [MSTWW]*Lemma A.2, we obtain that $T(R)$ is $p$-torsion free, as desired. ∎
@@ Appendix B Test ideal of graded ring
Since $S$ is quasi-Gorenstein, we obtain $\omega_{S}\simeq S(\nu)$ as graded $S$-modules, then $\nu<0$ by $a(S)<0$. By the isomorphism, the trace map $F_{*}\omega_{S}^{(p)}\to\omega_{S}$ is corresponding to $F_{*}(S(\nu))^{(p)}\to S(\nu)$, which is denoted by $u$. By the Matlis duality, it is enough to show that $\tau(S)=\bigcap_{e\in\mathbb{Z}_{\geq 0}}\mathrm{Im}(u^{e}\colon F^{e}_{*}S\to S)$, where $\tau(S)$ is the test ideal of $S$. First, we prove the following claim.
@@
Next, we prove the following claim.
@@
Next, we prove the following claim:
@@
Finally, we prove the assertion. We take rational numbers $01>t>\varepsilon>$ 0, then we have
@@
Appendix C Proof of Fedder-type criterion
@@
We use the notation introduced in Notation C.1 Let $n$ be a positive integer. Then $R$ is $n$-quasi-$F$-split if and only if there exists an $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that
@@
and in particular, $\psi$ satisfies conditions (a) and (b). Next, we assume that there exists an $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that $\psi$ satisfies conditions (a) and (b). Thus, the homomorphism $\psi$ induces the $\overline{R}$-module homomorphism $\overline{\psi}\colon Q_{R,n}\to\overline{R}$ such that $\overline{\psi}(1)$ is a unit. Replacing $\overline{\psi}$ by $\overline{\psi}(1)^{-1}\cdot\overline{\psi}$, we obtain a splitting of $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$, as desired. ∎
@@
We use the notation introduced in Notation C.1. Let $n$ be an integer with $n\geq 2$. We take $\varphi\in\operatorname{Hom}_{\overline{A}}(Q_{A,n},\overline{A})$ and $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$. By the isomorphism $Q_{A,n}\simeq F_{*}Q_{A,n-1}\oplus F_{*}\overline{A}$ in the proof of Theorem 3.7, we obtain the homomorphism $\psi\in\operatorname{Hom}_{\overline{A}}(F_{*}Q_{A,n-1},\overline{A})$ and $h\in\overline{A}$ such that
@@
for every $\alpha=(a_{0},\ldots,a_{n-1})\in Q_{A,n}$. Furthermore, we assume $\varphi$ satisfies condition (b) in Lemma C.2.
@@
We prove assertion (1) by induction on $n\geq 2$. We note that $h_{n}=g_{n}$ satisfies the desired condition for $s=n$. Since $\psi$ coincides with the decomposition
@@
and $\varphi$ satisfies condition (b) in Lemma C.2, the homomorphism $\psi$ induces the $\overline{R}$-module homomorphism $\psi_{R}\colon F_{*}Q_{R,n-1}\to\overline{R}$. Since $\overline{R}$ is Gorenstein, the homomorphism $\psi_{R}$ factors through the homomorphism induced by $u(F_{*}(f^{p-1}\cdot-))\colon F_{*}\overline{A}\to\overline{A}$. Therefore, there exists $\varphi^{\prime}\in\operatorname{Hom}_{\overline{A}}(Q_{A,n-1},\overline{A})$ such that $\psi$ and $u(F_{*}(f^{p-1}\cdot-))\circ\varphi^{\prime}$ induce same homomorphism $Q_{R,n}\to\overline{R}$ and $\varphi^{\prime}$ satisfies condition (b) in Lemma C.2. We take $g^{\prime}_{1},\ldots,g^{\prime}_{n-1}$ such that $\varphi^{\prime}=\varphi_{(g^{\prime}_{1},\ldots,g^{\prime}_{n-1})}$. By the isomorphism
@@
Next, we prove assertion (2). We take $r\geq j\geq 1$ and $a\in A$. Since $\varphi$ satisfies condition (b), we obtain
@@
where we define $\Delta_{0}(af_{j})=af_{j}$. For $n-s\geq 3$ or $p\neq 2$, we have $g_{s}\Delta_{n-s}(af_{j})\in I^{[p^{n+1-s}]}$ by 3.6 and assertion (1). Furthermore we have
@@
for every $n\geq 2$. Furthermore, $I^{\prime}_{n}\subseteq I^{\prime}_{n+1}$ for every $n\geq 1$ follows from the same argument. We prove assertion (1). First, we consider the following claim.
@@
Let $g\in A$ and $n\geq 1$. Then $g\in I_{n}$ if and only if there exists $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies condition (b) in Lemma C.2 and $g_{n}\equiv g\ \textrm{mod}\ I_{n-1}$, where if $n=1$, we regard $I_{0}$ as $(0)$.
@@
We prove the assertion by induction on $n$. For $n=1$, since $I_{1}=(I^{[p]}\colon I)$, $g\in I_{1}$ if and only if $u(F_{*}(g\cdot-))$ satisfies condition (b) in Lemma C.2, as desired. Next, we assume $n\geq 2$. We take $g\in I_{n}$, then there exists $g^{\prime}\in I_{n-1}$ such that $g-u(F_{*}(g^{\prime}\Delta_{1}(f^{p-1})))\in I_{1}$. By the induction hypothesis, there exist $g^{\prime}_{1},\ldots,g^{\prime}_{n-1}$ such that $\varphi_{(g_{1}^{\prime},\ldots,g^{\prime}_{n-1})}$ satisfies condition (b) in Lemma C.2, $g^{\prime}_{n-1}\equiv g^{\prime}\ \textrm{mod}\ I_{n-2}$, then by Lemma C.3, there exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g^{\prime}_{s}\equiv f^{p^{n-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n-s}]}$ for each $n-1\geq s\geq 1$. We define $\varphi^{\prime}$ by
@@
and $u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in I_{2}$, we obtain $g_{n}\equiv g\ \textrm{mod}\ I_{n}$. Thus, it is enough to show that $\varphi$ satisfies condition (b) in Lemma C.2. We take $g_{1},\ldots,g_{n-1}$ such that $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$, then we have
@@
for $\alpha\in W_{n-1}((f_{1},\ldots,f_{r}))$, we obtain that $\varphi$ satisfies condition (b) in Lemma C.2, as desired.
Next, we prove the converse direction, thus we take $g_{1},\ldots,g_{n}\in\overline{A}$ such that the corresponding homomorphism $\varphi:=\varphi_{(g_{1},\ldots,g_{n})}$ satisfies condition (b) in Lemma C.2, then it is enough to show that $g_{n}\in I_{n}$. For $n=1$, we have $g_{1}\in(I^{[p]}\colon I))=I_{1}$, as desired, thus we assume $n\geq 2$. By Lemma C.3, there exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g_{s}\equiv f^{p^{n+1-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n+1-s}]}$, and in particular, $h_{n}\equiv g_{n}\ \textrm{mod}\ I^{[p]}$. By the proof of Lemma C.3 and the induction hypothesis, we obtain that $h_{n-1}\in I_{n-1}$, thus we have $u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))\in I_{n}$. Furthermore, by the construction of $I_{2}$, we have
@@
Finally, we prove the assertion. We have
@@
Next, we prove assertion (2). We define the sequence of ideals $\{I^{e}_{n}\}$ by $I^{e}_{1}:=f^{p-1}u^{e}(F^{e}_{*}f^{p^{e}-1}\overline{A})+I^{[p]}$ and
@@
Let $g\in A$ and $n,e\geq 1$. Then $g\in I^{e}_{n}$ if and only if there exists $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies condition (b) in Lemma C.2, there exists $h\in f^{p^{e+n}-1}\overline{A}$ with $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, and $g_{n}\equiv g\ \textrm{mod}\ I^{e}_{n-1}$, where if $n=1$, we regard $I^{e}_{0}$ as $(0)$.
@@
For $n=1$, it follows from definition and the fact that condition (b) in Lemma C.2 is equivalent to $g_{1}\in f^{p-1}\overline{A}+I^{[p]}$. We assume $n\geq 2$. We take $g\in I^{e}_{n}$, then there exists $g^{\prime}\in I^{e}_{n-1}$ such that $g-u(F_{*}(g^{\prime}\Delta_{1}(f^{p-1})))\in I_{1}$. By induction hypothesis, there exist $g_{1}^{\prime},\ldots,g^{\prime}_{n-1}$ such that $\varphi_{(g^{\prime}_{1},\ldots,g^{\prime}_{n-1})}$ satisfies condition (b) in Lemma C.2, we have $u^{e}(F^{e}_{*}h^{\prime})=g^{\prime}_{1}$ for some $h^{\prime}\in f^{p^{e+n-1}-1}\overline{A}$, and $g^{\prime}_{n-1}\equiv g^{\prime}\ \textrm{mod}\ I^{e}_{n-2}$. By the proof of Claim C.4, if we put $g_{s}:=f^{p^{n-s}(p-1)}g^{\prime}_{s}$ for every $n-1\geq s\geq 1$ and
@@
then $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies condition (b) Lemma C.2 and $g\equiv g_{n}\ \textrm{mod}\ I^{\prime}_{n}$. We note that $I_{2}\subseteq I^{e}_{n}$ if $n\geq 3$. Furthermore, we have
@@
Next, we prove the converse direction, thus we take $g_{1},\ldots,g_{n}\in\overline{A}$ such that the corresponding homomorphism $\varphi:=\varphi_{(g_{1},\ldots,g_{n})}$ satisfies condition (b) in Lemma C.2 and there exists $h\in f^{p^{e+n}-1}\overline{A}$ such that $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, then it is enough to show that $g_{n}\in I^{e}_{n}$. For $n=1$, it follows from the definition of $I^{e}_{1}$. For $n\geq 2$, we take $h_{1},\ldots,h_{n}$ as in Lemma C.3, then the assertion follows from $h_{n-1}\in I^{e}_{n-1}$ by the induction hypothesis and
@@
for every $a\in\overline{A}$, and in particular, we have $u^{e}(F^{e}_{*}h)=g_{1}$. By Claim C.5, we have $g_{n}\in I^{e}_{n}=I^{\prime}_{n}$. Since $\varphi(1)=u(F_{*}g_{n})\notin\mathfrak{m}$, we have $I^{\prime}_{n}\nsubseteq\mathfrak{m}^{[p]}$, as desired. On the other hand, if $I^{e}_{n}=I^{\prime}_{n}\nsubseteq\mathfrak{m}^{[p]}$, then there exists $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$ such that $\varphi$ satisfies condition (b) in Lemma C.2, there exists $h\in f^{p^{e+n}-1}\overline{A}$ such that $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, and $u(F_{*}g_{n})\notin\mathfrak{m}$ by Claim C.5. Then $(\varphi,u^{e+n}(F^{e+n}_{*}h\cdot-))\colon Q_{A,n}\oplus F^{e+n}_{*}\overline{A}\to\overline{A}$ induces the homomorphism $Q_{R,n,e}\to\overline{R}$ and $(\varphi,u(F^{e}_{*}h\cdot-))(1)$ is a unit, as desired. ∎